20230629 UT



Night Info

  • Sunset: 19:37 MST
  • Evening 12-deg twilight: 20:35 MST
  • Night Center: 00:23 MST
  • Morning 12-deg twilight: 04:11 MST
  • Sunrise: 05:09 MST

Summary

See the google doc 20230629UT night log for details.

We started with MODS and observed GD153 (dual grating) and the SN2023ixf (DDT), followed by the photometric imaging field M13 in dual and direct modes. After that, Jenny and Paul G worked on the Shark NIR offsets until about midnight. The seeing had been degrading from the start and it was poor enough to preclude any LUCI2-AO work, so we reconfigured to the LBCs. It took some time to turn on the TMS lasers (Laser2 had not stabilized after 10 min and we rebooted the Etalon PC) and then, using the reference from the a few nights ago drove DX M1 into the negative Y limit and the initial images with the mirrors in this position were far from being copointed. We reset the mirrors to the collimation lookup table and collimated from scratch --- the initial DX pupil image was close to focus. I took a pair of extra/intra focal pupils to confirm that I needed to drop the Z global offset by ~0.8 mm. With this, we collimated and observed 2 SDSS+UKIDSS-Y fields & 1 Clem & Landolt UBVRI standard star field. I noted that, when collimated, the DX M1 Z positions were about 0.5mm below SX M1 Z positions. The seeing was still variable - from 1.2-1.7" - but we finished with a quick test of FPIA's Z22 correction and a short series of extra-focal pupils with TMS running.

Staff

  • OSA: J. Williams
  • ISA/SO: A.Becker
  • ECS Coordinator: J.Power
  • AO: G.Taylor & J.Power

Google doc spreadsheet with outline of plan

https://docs.google.com/spreadsheets/d/1JQxapZ0x6EmD8vzE1nmGI3I3eIBXEFnl/edit?usp=sharing&ouid=110700569789550059640&rtpof=true&sd=true

Tasks

SHARK NIR offset check (IT8911 investigations)

  • Check the rotator mode impacts to offsets
    • Rotator keywords IDLE and PARALLACTIC set for different presets with offsets applied
    • Send preset and offsets via IRC commands with the two different rotator keywords
    • Send presets and offsets via the SHARK clone with the two different rotator keywords
    • Monitor logs for IRC command output
  • IDL preset for VEGA
    • idl
    • IDL> iif_register, /TCS

      IDL> iif_preset, Dimm_3AlpLyr, MODE=ADAPTIVEACE_TRACK, RS=0

      IDL> iif_preset, /LEFT

      IDL>iif_runao

      IDL>iif_offset, (#X), (#Y)

DDT Target 2023ixf MODS

"A MODS spectrum would be great, especially if it's not too much work for you guys. The supernova is still at 11.5 mag in r band so it can be observed through clouds or at twilight if need be. Ideally we'd like a MODS spectrum taken at a lower airmass and at parallactic with an exposure time of 6x120sec (which is what I used for MMT Binospec and we haven't had any saturation issues). If Feige 34 could be taken as the standard that would be great, though as long as the standard is not HZ44 we should be fine." From Jeniveve Pearson 20230626
  • GD153 for standard
  • Scripts located in: /home/lbto/Service_Observing/2023A_D_E/2023A/MODS/M101_SN
  • Scripts tentatively generated for 4:15UT with PA139 but PA and slitwidth should be adjusted for time of observation as needed

LBC_Standard_Fields

  • Scripts located in: /home/lbto/Service_Observing/2023A_D_E/2023A/LBC/New_PhotFields
  • Fields
  • 20230626 only got 1 sdssY and 2 Clem & Landolt field so try to get 2 sdssY fields, either 2x on sdssY2 which is up all 2nd half - could use Y1 early and Y3 late.
  • Clem+Landolt fields: J175954+452015
  • Clem+Landolt fields:GD391
  • SDSS DR12: sdssY2_J2044d01 (2x)
  • Clem+Landolt fields: J205245+405115

MODS_Standards

  • Imaging Standard (priority 2)
    • M13 Field or M92
    • photometric conditions required
    • Scripts located in: /home/lbto/Service_Observing/2023A_D_E/2023A/MODS/Phot_Img
  • MODS Spectrophotmetric Standard
    • GD153 in Dual Grating (Priority 1)
      • Script for dual grating only located with SN2023ixf scripts
    • GD153 in all grating modes (priority 3)
      • photometric conditions required
      • Script that runs through all modes located in: /home/lbto/Service_Observing/2023A_D_E/2023A/MODS/SpectroPhotSTDS

LUCI2 AO imaging test (priority 1)

  • Juan Carlos PSF morphology (data was not fully obtained on UT20230626 because of technical issues)
    • Scripts and OT located in: /home/lbto/Service_Observing/2023A_D_E/2023A/LUCI/AO_PSFmorphology
      GDR3_132805941_wsky.xml generated for Target3 LUCI2 only
    • Optical loop data (TNx3) should be taken by AO operator for each on source image
  • PDS456 - LUCI2 AO imaging and LUCI1 Seeing limited, run through science filters H, Ks, J, z
    • Obtain a PSF calibrator
    • Scripts located in: /home/lbto/Service_Observing/2023A_D_E/2023A/LUCI/AO_and_ESM_testing/PDS456
    • They have been adjusted for LUCI1 Seeing limited:
      • PDS456_L2AO_L1SL.xml
      • PDS456_PSFStar_L2AO_L1SL.xml

LBC_Thermal_testing

This test will collect a series of extra-focal pupil images like FPIA takes (but with a specific OB and using all 4 chips) while driving the mirror temperatures up and down. In past tests, it has taken 6 hours to complete 1/2 sine curve. The pupil images will be analyzed to understand the aberrations that develop as a function of temperature gradients on the primaries. Seeing < 1.4" needed.
  • One circumpolar field that does not change elevation will be tracked.
    • const_elev_20h80d.xml
  • Readme file & scripts in ~/Service_Observing/2023A_D_E/2023A/LBC/Thermal_Pupil/

Night Plan

-- JenniferPower - 08 Jun 2023
Topic revision: r7 - 29 Jun 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback